# Objective and subjective evaluation of speech enhancement methods in the UDASE task of the 7th CHiME challenge

### v1

@@ 1 Introduction
The speech enhancement task consists of improving the quality and intelligibility of a degraded speech signal recording. One approach to achieve this is through noise suppression algorithms, which aim to estimate the clean speech signal by removing the additive background noise in the recording. Over the past 50 years, traditional signal-processing-based speech enhancement algorithms (Boll, 1979; Lim & Oppenheim, 1979; Ephraim & Malah, 1984; Martin, 2005; Loizou, 2013) have been progressively outperformed by data-driven approaches using hidden Markov models (HMMs) (Sameti et al., 1998; Ephraim, 1992; Sameti et al., 1998), codebook-based approaches (Srinivasan et al., 2005), non-negative matrix factorization (NMF) (Mohammadiha et al., 2013), and more recently deep learning (Wang & Chen, 2018).
@@
A large research effort has been put recently on the three main ingredients of supervised speech enhancement (the model, the metric, and the labeled dataset), and it is undeniable that this effort has led to unprecedented results, e.g., Weninger et al. (2015); Fu et al. (2017); Pascual et al. (2017); Choi et al. (2018); Zhao et al. (2018); Fu et al. (2019); Défossez et al. (2020); Cosentino et al. (2020); Hao et al. (2021); Pandey & Wang (2021); Fu et al. (2021); Richter et al. (2023). However, the fully supervised learning paradigm also has limitations when applied to speech enhancement. First, creating a synthetic dataset of realistic noisy speech mixtures is not easy and requires important engineering efforts. Second, supervised speech enhancement is effective as long as the acoustic recording conditions at test time are covered by the synthetic training data. This is a constraint that is hard to satisfy due to the variability of the acoustic recording conditions, in terms of noise type, signal-to-noise ratio, recording equipment, speaker-to-microphone distance and orientation, reverberation, etc. Supervised speech enhancement performance can thus decrease significantly in case of mismatch between the training and testing conditions (Pandey & Wang, 2020; Bie et al., 2022; Richter et al., 2023; Gonzalez et al., 2023). Finally, when the test domain differs from the synthetic training domain, supervised learning necessitates rebuilding the synthetic training dataset and retraining the model, which is time-consuming and computationally intensive. A more effective approach would be to automatically adapt models to real, unlabeled noisy speech recordings, eliminating the need for ground-truth clean speech signals.
@@
The ability to adapt to unknown adverse acoustic conditions while perceiving speech is a fundamental property of the human auditory system (Bent et al., 2009; Brandewie & Zahorik, 2010; Cooke et al., 2022), which cannot be reproduced in machine listening systems using a fully supervised approach. Adaptation of speech enhancement systems using real unlabeled noisy speech data is the main challenge that the CHiME-7 UDASE [^1] Unsupervised Domain Adaptation for conversational Speech Enhancement. task tried to address (Leglaive et al., 2023). Previous challenges for single-channel speech enhancement, such as the deep noise suppression (DNS) challenges (Reddy et al., 2020, 2021a, 2021b; Dubey et al., 2022), focused on a supervised setup with a training set consisting of a large amount of labeled synthetic data intended to cover diverse conditions. The CHiME-7 UDASE task was intended to study a different situation, targeting single-channel speech enhancement in a specific domain for which no well-matched labeled data are available for training.
@@
In Leglaive et al. (2023), we introduced the task and described the data and the baseline system. In the present paper, we extend the description of the CHiME-7 UDASE task by introducing the speech enhancement methods that were submitted to the challenge, describing their objective and subjective evaluation, and providing an analysis of the results. Along with this paper, we release the Javascript experimental platform we developed for the listening test, following the ITU-T P.835 (2003) methodology. We also release the audio files that were submitted to the challenge along with the corresponding human listening scores (in addition to various objective evaluation scores), which could serve as a voice quality dataset for speech enhancement research. [^2] https://doi.org/10.5281/zenodo.10418311 To the best of our knowledge, this is the first dataset of ITU-T P.835 (2003) subjective evaluation results made publicly available.
@@
The paper is organized as follows. The task and datasets are presented in Section 2. The objective and subjective evaluation protocols are described in Sections 3 and 4, respectively. Section 5 introduces the speech enhancement methods that participated in the CHiME-7 UDASE task. The results of the subjective and objective evaluations are presented and analyzed in Section 6 before concluding in Section 7.
@@ 2.1 Task
The CHiME-7 UDASE task focuses on single-channel speech enhancement in a specific target domain for which no well-matched labeled training data are available. It consists of using unlabeled data in the target domain to adapt supervised speech enhancement models trained on synthetic labeled data in a source domain. The target domain corresponds to the real conversational speech recordings of the CHiME-5 dataset (Barker et al., 2018). These recordings were made during dinner parties in real homes, with multiple speakers in noisy and reverberant environments. Given a mixture of one or more reverberant speakers and additive background noise, the objective of this task is to estimate the clean, potentially multi-speaker, reverberant speech, removing the additive background noise. The task is motivated by an assistive listening use case, in which a speech enhancement algorithm can help any individual to better engage in a conversation, by improving the overall speech quality and intelligibility within the ambient noise.
@@ 2.2 Datasets
    The CHiME-5 dataset (Barker et al., 2018) for the in-domain unlabeled data, which are used for training, development, and evaluation;
@@
    The LibriMix dataset (Cosentino et al., 2020) for the out-of-domain labeled data, which are used for training and development only;
@@
All three datasets include mixtures of reverberant speech and noise, with up to three overlapping speakers. All signals are sampled at 16 kHz. The datasets are briefly presented in the rest of this section, and additional information can be found in Leglaive et al. (2023).
@@ 2.3 CHiME-5 in-domain unlabeled data
The CHiME-5 dataset (Barker et al., 2018) consists of recordings made during twenty real dinner parties (or sessions) of between two and three hours. Each dinner party involved four participants wearing binaural microphones and took place in a different home, with three recording locations per home (kitchen, dining room, living room). The CHiME-5 recordings include natural conversations between multiple speakers in reverberant and noisy environments, and they are fully transcribed. Using the CHiME-5 transcription files, we estimated that 22% of the audio recordings contain only noise, 51% contain one single active speaker, and 20%, 5%, and 2% contain two, three, and four overlapping speakers, respectively. [^3] These numbers were obtained without considering any constraint on the overlap duration and using a labeling precision of 0.01 second. The CHiME-7 UDASE task only uses the binaural recordings (Kinect recordings are also available), from which single-channel audio segments with up to 3 overlapping speakers and background noise were extracted.
@@
The segmented CHiME-5 dataset used for the CHiME-7 UDASE task is summarized in Table 2.3. Additional details about the segmentation of the CHiME-5 original recordings are provided in Leglaive et al. (2023), and the scripts to generate the dataset are available online. [^5] https://github.com/UDASE-CHiME2023/CHiME-5
Figure 2: Instructions and rating scales for the ITU-T P.835 listening test.
@@
Subset # samples Mean STD ($H$:$M$:$S$)
@@
The design of the listening test for the CHiME-7 UDASE task followed the ITU-T P.835 (2003) methodology (Loizou, 2011). We were inspired by Hu & Loizou (2006, 2007); Naderi & Cutler (2021) who also evaluated speech enhancement algorithms following this methodology.
@@
Figure 2: Instructions and rating scales for the ITU-T P.835 listening test.
@@
Figure 5: Boxplots, violin plots, and mean results for the subjective BAK (4(a)), SIG (4(b)), and OVRL (4(c)) mean opinion scores of the ITU-T P.835 listening test. Black dots and numbers above the box/violin plots correspond to the mean results. In each figure, the systems are ranked according to their mean results.

### v2

@@ 1 Introduction
The speech enhancement task consists of improving the quality and intelligibility of a degraded speech signal recording. One approach to achieve this is through noise suppression algorithms, which aim to estimate the clean speech signal by removing the additive background noise in the recording. Over the past 50 years, traditional signal-processing-based speech enhancement algorithms [Boll, 1979, Lim & Oppenheim, 1979, Ephraim & Malah, 1984, Martin, 2005, Loizou, 2013] have been progressively outperformed by data-driven approaches using hidden Markov models (HMMs) [Sameti et al., 1998, Ephraim, 1992, Sameti et al., 1998], codebook-based approaches [Srinivasan et al., 2005], non-negative matrix factorization (NMF) [Mohammadiha et al., 2013], and more recently deep learning [Wang & Chen, 2018].
@@
A large research effort has been put recently on the three main ingredients of supervised speech enhancement (the model, the metric, and the labeled dataset), and it is undeniable that this effort has led to unprecedented results, e.g., Weninger et al. [2015], Fu et al. [2017], Pascual et al. [2017], Choi et al. [2018], Zhao et al. [2018], Fu et al. [2019], Défossez et al. [2020], Cosentino et al. [2020], Hao et al. [2021], Pandey & Wang [2021], Fu et al. [2021], Richter et al. [2023]. However, the fully supervised learning paradigm also has limitations when applied to speech enhancement. First, creating a synthetic dataset of realistic noisy speech mixtures is not easy and requires important engineering efforts. Second, supervised speech enhancement is effective as long as the acoustic recording conditions at test time are covered by the synthetic training data. This is a constraint that is hard to satisfy due to the variability of the acoustic recording conditions, in terms of noise type, signal-to-noise ratio, recording equipment, speaker-to-microphone distance and orientation, reverberation, etc. Supervised speech enhancement performance can thus decrease significantly in case of mismatch between the training and testing conditions [Pandey & Wang, 2020, Bie et al., 2022, Richter et al., 2023, Gonzalez et al., 2023]. Finally, when the test domain differs from the synthetic training domain, supervised learning necessitates rebuilding the synthetic training dataset and retraining the model, which is time-consuming and computationally intensive. A more effective approach would be to automatically adapt models to real, unlabeled noisy speech recordings, eliminating the need for ground-truth clean speech signals.
@@
The ability to adapt to unknown adverse acoustic conditions while perceiving speech is a fundamental property of the human auditory system [Bent et al., 2009, Brandewie & Zahorik, 2010, Cooke et al., 2022], which cannot be reproduced in machine listening systems using a fully supervised approach. Adaptation of speech enhancement systems using real unlabeled noisy speech data is the main challenge that the CHiME-7 UDASE [^1] Unsupervised Domain Adaptation for conversational Speech Enhancement. task tried to address [Leglaive et al., 2023]. Previous challenges for single-channel speech enhancement, such as the deep noise suppression (DNS) challenges [Reddy et al., 2020, 2021a, 2021b, Dubey et al., 2022], focused on a supervised setup with a training set consisting of a large amount of labeled synthetic data intended to cover diverse conditions. The CHiME-7 UDASE task was intended to study a different situation, targeting single-channel speech enhancement in a specific domain for which no well-matched labeled data are available for training.
@@
In Leglaive et al. [2023], we introduced the task and described the data and the baseline system. In the present paper, we extend the description of the CHiME-7 UDASE task by introducing the speech enhancement methods that were submitted to the challenge, describing their objective and subjective evaluation, and providing an analysis of the results. Along with this paper, we release the JavaScript experimental platform we developed for the listening test, following the ITU-T P.835 [2003] methodology. We also release the audio files that were submitted to the challenge along with the corresponding human listening scores (in addition to various objective evaluation scores), which could serve as a voice quality dataset for speech enhancement research [Leglaive et al., 2024]. To the best of our knowledge, this is the first dataset of ITU-T P.835 [2003] subjective evaluation results made publicly available.
@@
The paper is organized as follows. The task and datasets are presented in Section 2. The objective and subjective evaluation protocols are described in Sections LABEL:sec:objective_eval and LABEL:sec:listening_test, respectively. Section LABEL:sec:SE_methods introduces the speech enhancement methods that participated in the CHiME-7 UDASE task. The results of the subjective and objective evaluations are presented and analyzed in Section LABEL:sec:results before concluding in Section 7.
@@ 2.1 Task
The CHiME-7 UDASE task focuses on single-channel speech enhancement in a specific target domain for which no well-matched labeled training data are available. It consists of using unlabeled data in the target domain to adapt supervised speech enhancement models trained on synthetic labeled data in a source domain. The target domain corresponds to the real conversational speech recordings of the CHiME-5 dataset [Barker et al., 2018]. These recordings were made during dinner parties in real homes, with multiple speakers in noisy and reverberant environments. Given a mixture of one or more reverberant speakers and additive background noise, the objective of this task is to estimate the clean, potentially multi-speaker, reverberant speech, removing the additive background noise. The task is motivated by an assistive listening use case, in which a speech enhancement algorithm can help any individual to better engage in a conversation, by improving the overall speech quality and intelligibility within the ambient noise.
@@ 2.2 Datasets
    The CHiME-5 dataset [Barker et al., 2018] for the in-domain unlabeled data, which are used for training, development, and evaluation;
@@
    The LibriMix dataset [Cosentino et al., 2020] for the out-of-domain labeled data, which are used for training and development only;
@@
All three datasets include mixtures of reverberant speech and noise, with up to three overlapping speakers. All signals are sampled at 16 kHz. The datasets are presented in the rest of this section, and additional information can be found in Leglaive et al. [2023].
@@ 2.3 CHiME-5 in-domain unlabeled data
The CHiME-5 dataset [Barker et al., 2018] consists of recordings made during twenty real dinner parties (or sessions) of between two and three hours. Each dinner party involved four participants wearing binaural microphones and took place in a different home, with three recording locations per home (kitchen, dining room, living room). The CHiME-5 recordings include natural conversations between multiple speakers in reverberant and noisy environments, and they are fully transcribed. Using the CHiME-5 transcription files, we estimated that 22% of the audio recordings contain only noise, 51% contain one single active speaker, and 20%, 5%, and 2% contain two, three, and four overlapping speakers, respectively. [^2] These numbers were obtained without considering any constraint on the overlap duration and using a labeling precision of 0.01 second. The CHiME-7 UDASE task only uses the binaural recordings (Kinect recordings are also available), from which single-channel audio segments with up to 3 overlapping speakers and background noise were extracted.
@@
The segmented CHiME-5 dataset used for the CHiME-7 UDASE task is summarized in Table 2.3. Additional details about the segmentation of the CHiME-5 original recordings are provided in Leglaive et al. [2023], and the scripts to generate the dataset are available online. [^4] https://github.com/UDASE-CHiME2023/CHiME-5 (last accessed: 2024-02-02)
@@
Subset # samples Mean SD (HH:MM:SS)
@@
2.4 LibriMix out-of-domain labeled data
The CHiME-7 UDASE task employs the LibriMix dataset [Cosentino et al., 2020] for supervised learning on out-of-domain data. LibriMix was originally developed for speech separation in noisy environments, and it is derived from LibriSpeech “clean” utterances [Panayotov et al., 2015] and WHAM! noises [Wichern et al., 2019]. The Libri2Mix and Libri3Mix versions of the dataset contain noisy speech mixtures with 2 and 3 overlapping speakers, respectively. A single-speaker version of LibriMix (Libri1Mix) can be obtained by simply discarding one of the two speakers in Libri2Mix mixtures. A complete description of LibriMix is provided by Cosentino et al. [2020].
2.5 Reverberant LibriCHiME-5 close-to-in-domain labeled data
8pt. RT60 (s)
Subset Home Room # RIRs Mean SD
dev 2 Living room 160 0.582 0.023
2 Kitchen 120 0.500 0.022
2 Bedroom 136 0.471 0.056
3 Living room 136 0.437 0.026
3 Bedroom 120 0.431 0.036
eval 3 Kitchen 128 0.451 0.037
4 Living room 144 0.460 0.034
4 Kitchen 128 0.375 0.035
4 Bedroom 128 0.387 0.022
8pt.
Table 2: Estimated RT60s (in seconds) of the rooms in the VoiceHome corpus [Bertin et al., 2016] that were used to create the reverberant LibriCHiME-5 data (dev and eval subsets). The column ‘# RIRs’ gives the number of single-channel RIRs used to compute the mean and standard deviation (SD) of the RT60.
We created the reverberant LibriCHiME-5 dataset to provide “close-to-in-domain” labeled data for computing standard intrusive objective performance metrics used in speech enhancement, such as the scale-invariant signal-to-distortion ratio (SI-SDR) [Le Roux et al., 2019]. This dataset consists of synthetic mixtures of reverberant speech and noise, with up to three simultaneously active speakers, labeled with the clean reference speech signals. In-domain noise signals were extracted from the CHiME-5 recordings using the ground-truth transcriptions (see Section 2.3), and clean speech utterances were taken from the LibriSpeech dataset [Panayotov et al., 2015] and convolved with room impulse responses (RIRs) from the VoiceHome corpus [Bertin et al., 2016]. These RIRs were recorded in 12 different rooms of 3 real homes, with 4 rooms per home: living room (room 1), kitchen (room 2), bedroom (room 3), and bathroom (room 4). Bathrooms were excluded for the reverberant LibriCHiME-5 dataset. In each room, RIRs were recorded for 2 different positions and geometries of an 8-channel microphone array and 7 to 9 different positions of the loudspeaker. Table 2.5 indicates the estimated reverberation times (RT60s) of the rooms in the VoiceHome corpus whose RIRs were used to create the reverberant LibriCHiME-5 dataset. The RT60 is defined as the time it takes for the sound energy to decrease by 60 dB after the extinction of the source. For each room in each home, we estimated the RT60 on each single-channel RIR by fitting a straight line on Schroeder’s integrated energy decay curve using linear least-squares regression [Schroeder, 1965]. We then computed the mean RT60 and the standard deviation for each room.
The process to create the synthetic reverberant LibriCHiME-5 dataset was the following. For each mixture in the dataset, we randomly chose the maximum number $n\in\{1,2,3\}$ of simultaneously active speakers in the mixture, with $p(n=i)=0.60,0.35,0.05$ for $i=1,2,3$, respectively, which is consistent with the distribution of the segmented CHiME-5 dataset. Each mixture’s speakers were randomly sampled from the list of LibriSpeech speakers with equal probability to be a male or a female. We used the VoiceHome corpus to simulate the acoustic recording environment in the reverberant LibriCHiME-5 dataset. For each mixture, we randomly and successively sampled a home, a room, an array position/geometry, $n$ speaker positions without replacement, and a channel of the microphone array, which gave a set of RIRs. LibriSpeech utterances were convolved with the selected RIRs to obtain the reverberant speech utterances. These were then mixed following speech activity patterns extracted from the CHiME-5 transcription files (diarization labels) to simulate a natural conversation between multiple speakers. This synthetic mixing of multi-speaker speech involved selecting and trimming LibriSpeech utterances in order to make them fit in the CHiME-5 activity patterns. The multi-speaker reverberant speech and noise mixtures were finally created such that the per-speaker signal-to-noise ratio (SNR) was distributed as a Gaussian with a mean of 5 dB and a standard deviation of 7 dB to match the SNR distribution of the CHiME-5 dataset as estimated by Brouhaha [Lavechin et al., 2023]. This was achieved by first sampling a global per-mixture SNR $x\sim\mathcal{N}(5,\sigma_{1}^{2})$ and then sampling a local per-speaker SNR $y_{n}\sim\mathcal{N}(x,\sigma_{2}^{2})$, with $\sigma_{1}=6.7082$ and $\sigma_{2}=2$ ($\sqrt{\sigma_{1}^{2}+\sigma_{2}^{2}}\approx 7$ dB). The value of $\sigma_{2}$ was chosen such that the loudness difference between multiple speakers remained moderate; this was again to simulate a conversation. A detailed description of the process implemented to create the reverberant LibriCHiME-5 mixtures (metadata and audio files) and the corresponding Python code are available online. [^5] https://github.com/UDASE-CHiME2023/reverberant-LibriCHiME-5 (last accessed: 2024-04-03).
Overall, in the reverberant LibriCHiME-5 dataset, the speech utterances were convolved with RIRs measured in real homes, the noise signals were extracted from the in-domain CHiME-5 recordings, the SNR was chosen to approximately match that of the target domain, and the speech utterances were mixed to simulate a conversation using the CHiME-5 transcription. It is therefore hoped that the performance on the reverberant LibriCHiME-5 dataset corresponds to an estimate of the performance on the CHiME-5 dataset, which will be confirmed by the analysis of the evaluation results. The reverberant LibriCHiME-5 dataset is summarized in Table LABEL:tab:librichime-5.
(b) SIG MOS results.
(c) OVRL MOS results.
@@ 2.5 Reverberant LibriCHiME-5 close-to-in-domain labeled data
Figure 4: Boxplots, violin plots, and mean results for the subjective BAK (LABEL:fig:listening_test_BAK), SIG (4(b)), and OVRL (4(c)) mean opinion scores of the ITU-T P.835 listening test. Black dots and numbers above the box/violin plots correspond to the mean results. In each figure, the systems are ranked according to their mean results.
6.2.5 Comparison with the objective evaluation
Figure 5: Pearson correlation between objective and subjective evaluation results, computed using only single-speaker samples (115 samples in total).
It can be interesting to compare the results of the objective and subjective evaluations. The CMGAN submissions, which were found to perform best in terms of DNSMOS metrics on both CHiME-5 and reverberant LibriCHiME-5 datasets, are the worst-performing systems according to the listening test (see Figure 2.3). As detailed in Close et al. [2023], this is probably due to the training of the CMGAN models that rely on the optimization of the supervised DNSMOS metrics. It appears that the models learned to maximize the metrics while providing output speech signals of unsatisfactory quality. This is probably because the DNSMOS model is supervised and it can thus provide unexpected or incoherent results when fed with signals that are too far from its training conditions. On the contrary, the objective evaluation results using standard objective performance metrics (SI-SDR, PESQ, and STOI) computed on the close-to-in-domain reverberant LibriCHiME-5 dataset are consistent with the subjective evaluation results on the CHiME-5 dataset, in terms of ranking of the methods. This again confirms that reverberant LibriCHiME-5 can be used to approximate in-domain performance in the context of the CHiME-7 UDASE task.
Finally, Figure 5 shows the PCC between the subjective SIG, BAK, OVRL MOS, and the nonintrusive DNSMOS and TorchAudio-Squim metrics. It can be seen that, apart from a PCC of 0.73 between the DNSMOS BAK score and the subjective BAK MOS, the nonintrusive objective metrics poorly correlate with the subjective ones (PCC lower than or equal to 0.50). The DNSMOS BAK score was also found to be the most reliable nonintrusive metric when analyzing the objective evaluation results in Section LABEL:subsec:obj_eval. Interestingly, the correlation between the subjective SIG, BAK, and OVRL metrics (0.17, 0.76, 0.66) is very close to the correlation between the equivalent objective metrics as provided by DNSMOS (0.18, 0.69, 0.66). This suggests that DNSMOS has appropriately learned the correlation between the SIG, BAK, and OVRL scores, but unfortunately, this is not sufficient to ensure good generalization, in particular on the CHiME-7 UDASE data.
7 Conclusion
Fully-supervised speech enhancement models are trained – and most of the time also evaluated – using only synthetic data, which cannot always capture the diversity of real-world speech recordings. A strong mismatch between the synthetic training domain and the real test domain will significantly affect the performance of the model. To address this issue, the CHiME-7 UDASE task aimed to leverage real-world unlabeled noisy speech recordings from the test domain for the unsupervised adaptation of speech enhancement models. Evaluating unsupervised domain adaptation methods for speech enhancement is by definition a challenging task because the ground-truth clean speech signals in the target domain are not available. In this paper, we presented the methodology and analyzed the results of the objective and subjective evaluations conducted in the CHiME-7 UDASE task.
Objective evaluation in the target domain (defined by the CHiME-5 recordings) relied on the recent DNSMOS P.835 and TorchAudio-Squim nonintrusive performance metrics. Complementarily, we developed the synthetic labeled reverberant LibriCHiME-5 dataset for objective evaluation with common intrusive metrics (SI-SDR, PESQ, STOI) on close-to-in-domain data. The subjective evaluation consisted of an ITU-T P.835 listening test, which was implemented in a JavaScript experimental platform and conducted in person at the University of Sheffield (UK). The subjective evaluation revealed the difficulty of the CHiME-7 UDASE task. Among the four speech enhancement systems that were evaluated in the listening test, only one succeeded in improving the overall quality compared to the unprocessed noisy speech. All systems successfully reduced the background noise but always at the expense of increased distortion, which eventually affected the overall perceived quality.
The analysis of the results revealed that the DNSMOS P.835 and TorchAudio-Squim nonintrusive performance metrics should be used with caution. Their effectiveness was demonstrated for the specific benchmarks presented in Reddy et al. [2022] and Kumar et al. [2023]. However, except for the DNSMOS BAK score, these metrics were shown to poorly correlate with the subjective ratings. On the contrary, the ranking of speech enhancement methods on reverberant LibriCHiME-5 using more traditional intrusive objective performance metrics was similar to the ranking based on subjective evaluation. This shows that the reverberant LibriCHiME-5 dataset can be used for in-domain evaluation of speech enhancement models adapted to the unlabeled CHiME-5 dataset. While this is useful practically, it is not entirely satisfying as unsupervised domain adaptation methods should in principle be evaluated on the unlabeled data of the target domain, which is very challenging methodologically [You et al., 2019, Musgrave et al., 2021]. Further research is needed to address the challenge of training and evaluation of speech enhancement models without clean speech labels.
Appendix A Experimental platform for the listening test
Screenshots of the JavaScript experimental platform developed for the listening test of the CHiME-7 UDASE task are shown in Figure 2.3.
(a) SIG rating scale and instruction.
(b) BAK rating scale and instruction.
(c) OVRL rating scale and instruction.
Figure 6: Windows of the experimental platform used for the listening test. Each window presents an instruction and a rating scale following the ITU-T P.835 recommendation, and a 5-position slider to register the vote.
Acknowledgments
The authors thank the CHiME Steering Group for their support in the organization of the CHiME-7 UDASE task and the participants of the listening test organized at the University of Sheffield. S. Leglaive acknowledges support from the French National Research Agency (ANR) as a part of the DEGREASE project (ANR-23-CE23-0009).